OpenStudy (anonymous):

4 log 2 - log 3 + log 16 . rewrite the expression as a single log

OpenStudy (anonymous):

log[(16*16)/3= log (256/3)

OpenStudy (anonymous):

= log (2^4 * 16) / 3 = log (256/3)